At the beginning of a​ walk, Roberto and Juanita are 5.5

miles apart. If they leave at the same time and walk in the same​ direction, Roberto overtakes Juanita in 11 hours. If they walk towards each​ other, they meet in 1 hour. What are their​ speeds?

Answer by ankor@dixie-net.com(22740)

At the beginning of a​ walk, Roberto and Juanita are 5.5 miles apart.
If they leave at the same time and walk in the same​ direction, Roberto overtakes Juanita in 11 hours.
If they walk towards each​ other, they meet in 1 hour.
What are their​ speeds?
:
Let r = R's speed
let j = J's speed
:
Write a distance equation for each scenario. dist = time * speed
11(r-j) = 5.5, walking in the same direction
1(r+j) = 5.5, walking towards each other
:
simplify the 1st equation, divide both sides by 11, add to the 2nd equation
r - j = .5
r + j = 5.5
-------------addition eliminates j, find r
2r + 0 = 6.0
r = 6/2
r = 3 mph is R's speed
Find j
3 + j = 5.5
j = 5.5- 3
j = 2.5 mph is J's speed
:
;
Check this in first original equation
11(3 - 2.5) =
11(.5) = 5.5